OpenStudy (anonymous):

i need help finding the x and y intercepts of a rational function PLEASE HELP!!!1

OpenStudy (anonymous):

and the rational function is...?

OpenStudy (anonymous):

(x^2+x-2)/x^2-3x-4)

OpenStudy (anonymous):

i know that when i factor this out it is (x+2)9x-1)/(x+1)(x-4)

OpenStudy (anonymous):

i mean (x+2)(x-1)/(x+1)(x-4)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so what do you think would be the values of x that would make your denominator equal to 0?

OpenStudy (anonymous):

^this is to determine the limits of the x

OpenStudy (anonymous):

are they (1,0) (-2,0) and (0, 0.5)

OpenStudy (anonymous):

where did 0.5 come from?

OpenStudy (anonymous):

wouln't that be they y intercept?

OpenStudy (anonymous):

oh yeah sorry...i was imagining x

OpenStudy (anonymous):

ok how would i find the domain for this function?

OpenStudy (anonymous):

the domain...as i said a while ago...is the limit of x...so find the values of x that would make the denominator = 0

OpenStudy (anonymous):

would it be \[(\infty, -1) \cup (-1,4) \cup (4, \infty)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

except the first should be (\(-\infty\), -1)

OpenStudy (anonymous):

yes my mistake another question is how can i find the horizontal asymptote

OpenStudy (anonymous):

horizontal asymptote is the value that makes x equal to infinity

OpenStudy (anonymous):

value of y^

OpenStudy (anonymous):

ok so does that mean that there is no horizontal aymptote in this function?

OpenStudy (anonymous):

do you know how to solve for the range?

OpenStudy (anonymous):

you have to solve the range first before judging

OpenStudy (anonymous):

the range is all real numbers

OpenStudy (anonymous):

so does that mean that the horizontal asymptote is none

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

If you have a rational function where the numerator and denominator have the same degree, which they do here, the asymptote can be found by dividing the coefficients of the highest degree term. So, the numerator has a x^2 with coefficient 1, and the denominator has a x^2 with coefficient 1, so the asymptote is y=1/1=1

OpenStudy (anonymous):

does that mean that the horizontal asymptote is 1?

OpenStudy (anonymous):

but i thought that the horizontal asymptote was 0

OpenStudy (anonymous):

i just need to know what the horizontal asymptote is

Directrix (directrix):

>does that mean that the horizontal asymptote is 1? @redjess13 Yes, the horizontal asymptote is y = 1.